OpenStudy (anonymous):

Which sequence is modeled by the graph below? coordinate plane showing the points 1, 1; 2, 2; 3, 4; and 4, 8 an = 1(2)n − 1 an = 2(1)n − 1 an = 1 + (2)n − 1 an = (2)(n − 1) @ganeshie8

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Jhannybean plz help

OpenStudy (anonymous):

i would choose #4, if I understand the notation correctly. An is equal 2 to the power of (n-1)

OpenStudy (perl):

i agree , if a(n) = 2^(n-1)

OpenStudy (anonymous):

thank you very much :) @perl @Sergii

OpenStudy (anonymous):

@perl i'm confused @sergii said that it was #4 and u said that its 2^(n-1) and thats 1 and 2

OpenStudy (perl):

can you upload the original question, its a bit difficult to read

OpenStudy (anonymous):

Which sequence is modeled by the graph below? coordinate plane showing the points 1, 1; 2, 2; 3, 4; and 4, 8 an = 1(2)n − 1 an = 2(1)n − 1 an = 1 + (2)n − 1 an = (2)(n − 1)

OpenStudy (perl):

your choices are difficult to understand

OpenStudy (perl):

can you take a screen shot of it

OpenStudy (anonymous):

@perl

OpenStudy (perl):

ok then we want this function a(n)= 1*2^(n-1)

OpenStudy (anonymous):

so it would be A right?

OpenStudy (perl):

the choices are : a) an = 1*2^(n-1) b) an = 2*1^(n-1) c) an = 1 + 2^(n-1) d) an = (2)*(n-1)

OpenStudy (perl):

yes that is what I get

OpenStudy (mathmath333):

\(\Large 1, 1;~~ 2, 2;~~ 3, 4;~~ 4, 8\\~\\ \implies \Large 1, 2^{1-1};~~ 2, 2^{2-1};~~ 3, 2^{3-1};~~ 4, 2^{4-1|}\\~\\\)

OpenStudy (anonymous):

Hentai is the Answer

OpenStudy (anonymous):

Always the Answer :)

OpenStudy (anonymous):

I am with @mathmath333 Although both A and D are correct but D is precisely correct.

OpenStudy (anonymous):

Its hentai

OpenStudy (anonymous):

YOUR WRONG IT's HENTAI

OpenStudy (anonymous):

HENTAI